# Deterministic Dynamic Edge-Colouring

### v1

@@ Abstract
Given a dynamic graph $G$ with $n$ vertices and $m$ edges subject to insertion an deletions of edges, we show how to maintain a $(1+\varepsilon)\Delta$-edge-colouring of $G$ without the use of randomisation. More specifically, we show a deterministic dynamic algorithm with an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ using $(1+\varepsilon)\Delta$ colours. If $\varepsilon^{-1}\in 2^{O(\log^{0.49}n)}$, then our update time is sub-polynomial in $n$.
While there exists randomised algorithms maintaining colourings with the same number of colours [Christiansen STOC’23, Duan, He, Zhang SODA’19, Bhattacarya, Costa, Panski, Solomon SODA’24] in polylogarithmic and even constant update time, this is the first deterministic algorithm to go below the greedy threshold of $12\Delta-$ 1 colours for all input graphs. On the way to our main result, we show how to dynamically maintain a shallow hierarchy of degree-splitters with both recourse and update time in $n^{o(1)}$. We believe that this algorithm might be of independent interest.
@@ 1 Introduction & Related Work
Edge-colouring is a classic and fundamental problem studied extensively in both graph theory as well as many different models of computation [${}^{+}$85, Ber22, Chr23, DHZ19, BGW21]. Given a simple graph $G=(V,E)$, a $k$-edge-colouring of $G$ is a function $c:E(G)\rightarrow[k]$ that maps every edge to a colour in some colour set $[k]=\{1,2,\dots,k\}$, An edge-colouring is proper if $e\cap e^{\prime}\neq\emptyset$ implies that $c(e)\neq c(e^{\prime})$ for any pair of distinct edges $e,e^{\prime}$. The chromatic index $\chi^{\prime}(G)$ of $G$ is then the smallest $k$ for which a proper $k$-edge-colouring of $G$ exists. It is not difficult to check that $\chi^{\prime}(G)\geq\Delta$ if $G$ has maximum degree $\Delta$. A classic result of Vizing [Viz64] shows that in fact $\chi^{\prime}(G)\leq\Delta+1$. Deciding whether $\chi^{\prime}(G)=\Delta$ or $\chi^{\prime}(G)=\Delta+1$ is, however, not easy, and it was shown by Holyer that determining $\chi^{\prime}(G)$ exactly is, in fact, NP-complete [Hol81].
@@
Recently, the edge-colouring problem has received a lot of attention in the dynamic algorithms community. Here one studies dynamic graphs, where edges are continually inserted and deleted, and the challenge is to maintain some proper edge-colouring of the graph as it is subjected to updates. There has been some exciting progress on developing efficient randomised dynamic algorithms using only $(1+\varepsilon)\Delta$ colours. This number of colours go significantly below the greedy-threshold of $12\Delta-$ 1 colours. Duan, He, and Zhange [DHZ19] showed how to maintain such a $(1+\varepsilon)\Delta$-edge-colouring with an update time of $O(\tfrac{\log^{7}n}{\varepsilon^{2}})$ provided that $\Delta=\Omega(\tfrac{\log^{2}n}{\varepsilon^{2}})$. Christiansen got rid of all requirements on $\Delta$ while maintaining a worst-case update time in $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$. Very recently Bhattacarya, Costa, Panski, and Solomon [BCPS] pushed the update time all the way down to $\tilde{O}(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ under some additional assumptions on $\Delta$ and $\varepsilon$.
@@
However, all of the above algorithms rely on randomness in several and crucial ways, and the landscape of deterministic and dynamic algorithms is a lot less colourful. Barenboim and Maimon [BM17] gave a dynamic algorithm with $O(\sqrt{\Delta})$ worst-case update time using $O(\Delta)$ colours. This was subsequently improved by Bhattacarya, Chakrabarty, Henzinger and Nanongkai [BCHN18] who gave a fully dynamic algorithm for computing a $(2\Delta-1)$-edge-colouring with worst case $O(\log\Delta)$ update time. Bhattacarya, Costa, Panski, and Solomon [BCPS23a] and independently Christiansen, Rotenberg, and Vlieghe [CRV23] showed how to maintain a $(\Delta+O(\alpha))$-edge-colouring in amortised $O(\operatorname{\text{{\rm poly}}}(\log n))$ update-time, where $\alpha$ is the arboricity of $G$. However, if $\Delta\sim\alpha$, this algorithm will not always use a number of colours below the greedy threshold as $\alpha$ can be as big as $\tfrac{\Delta}{2}$. In particular, until now no deterministic and dynamic algorithm always using less than $12\Delta-$ 1 colours with sub-linear update time was known.
@@
Theorem 1 (Informal).
Let $G$ be a dynamic graph and $\varepsilon^{-1}\leq 2^{\operatorname{\text{{\rm poly}}}(\log\log n)}$ a given parameter. Then, we can maintain a proper $((1+\varepsilon)\Delta)$-edge-colouring of $G$ in $n^{o(1)}$ amortised update time per operation.
@@ 1.1 Main technical challenges
In this section, we will pinpoint two issues towards designing deterministic dynamic edge-colouring algorithms with few colours. These are barriers we know how to overcome in either the static setting or by using randomisation, but so far, to the best of our knowledge, no current techniques deal with these issues directly.
@@
Many static algorithms, distributed algorithms, and dynamic randomised algorithms produce edge-colourings with few colours via a two-step approach. The first step is a form for dimenstionality reduction; one attempts to reduce the problem to a case where $\Delta$ is small. The second step is then to apply an algorithm that is fast and efficient for graphs with small maximum degree.
@@
Many static algorithms [${}^{+}$85, BCPS23b, Sin19] and many distributed algorithms [GKMU18, ${}^{+}$20, GS17] adress the first step by computing a hierarchy of degree-splitters. A degree-splitter splits a graph $G$ into two covering subgraphs $G_{1},G_{2}$ such that the maximum degree of $G_{i}$ is roughly half of the maximum degree of $G$. By applying this procedure recursively one can reduce the problem to $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ sub-problems each of maximum-degree $\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)$. However, these static and precise approaches for splitting the degree, seems less compatible with changes to the graph.
@@
If the maximum degree is large enough, it is easy to realise something similar in the randomised setting, even dynamically against an oblivious adversary. Indeed, one can place each edge into one of $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ subgraphs uniformly at random. It then follows from a standard application of Chernoff-bounds that the maximum degree of each such subgraph is split nicely. A different approach for reducing $\Delta$ is to sample uniformly at random a palette $S\subset[t]$ of colours such that the graph induced by all edges coloured with a colour from $S$, say $G[S]$, exhibits useful properties. Indeed, one can show that with high-probability the maximum degree of $G[S]$ is small and every vertex in $G$ has a free colour in $S$. One can then colour a new edge, by viewing it as being inserted into $G[S]$. This is the approach taken in [Chr23, DHZ19]. It is, however, not clear how to perform such a reduction in the deterministic dynamic setting; naively building a hierarchy can easily result in an insurmountable recourse (see section 2), and sampling colours or edges relies crucially on the power of randomisation.
@@
The second challenge is to efficiently colour a low-degree graph. In the static setting one could rely on parallelism [${}^{+}$85], and in the dynamic & distributed settings recent progress has shown that the recourse for extending colourings on low-degree graphs is small [Ber22, BD23, Chr23, DHZ19] and, more importantly, can be efficiently realised using randomisation [Ber22, BD23, Chr23, DHZ19]. However, naively searching for such an extending sequence of recolourings can be expensive, and so it is unclear how to locate it without using randomisation.
@@ 1.2 Further Related Work
Gabow, Nishizeki, Kariv, Leven, and Tereda [${}^{+}$85] showed how to compute a $(\Delta+1)$-edge-colouring in $\tilde{O}(m\sqrt{n})$ or $\tilde{O}(m\Delta)$ time. This improved over the $O(mn)$ time algorithm implied by the work of Vizing [Viz64]. Recently this was improved by Sinnamon [Sin19], and in special settings by Bernshteyn and Dhawan [BD23] and by Bhattacarya, Costa, Panski, and Solomon [BCPS23b]. Alon [Alo03] and Cole, Ost and Schirra [COS01] gave a near linear-time algorithm for $\Delta$-edge-colouring bipartite graphs. By applying a reduction due to Karloff and Shmoys [KS87], these algorithms yield $3\lceil\frac{\Delta}{2}\rceil$-edge-colouring algorithms for general graphs. When $\Delta=\Omega{(\varepsilon^{-1}\log n)}$, Duan, He and Zhang [DHZ19] gave a $O(\varepsilon^{-2}m\log^{6}n)$ time algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring. This has recently been improved by Bhattacarya, Costa, Panski, and Solomon [BCPS] to $\tilde{O}(m\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ time for many values of $\Delta$ and $\varepsilon$.
The edge-colouring problem has also been extensively studied in the LOCAL model [Ber22, BD23, ${}^{+}$18, Chr23, GKMU18, ${}^{+}$20, GS17, SV19] (see for instance [${}^{+}$18, GKMU18] for extensive surveys). Some of these algorithms rely on degree-splitting either rely on or directly consider the degree-splitting problem [GKMU18, ${}^{+}$20, GS17].
@@ 1.3 Results
Theorem 2.
Let $G$ be a dynamic graph, and $10<\varepsilon<$ 1 a given parameter. We can maintain a proper $(1+\varepsilon)\Delta$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised [^1] Here $\tilde{O}_{\log\varepsilon^{-1}}$ hides $\log\log n$ and $\log\varepsilon^{-1}$ factors. update time per operation.
@@
Theorem 3.
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ through the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-splitter of $G$ with an amortised update-time of $O(t^{3}\varepsilon^{-4}\log^{3}m)$.
@@
Theorem 4.
Let $G$ be a dynamic graph, and let $01>\varepsilon>$ 0 and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
@@
  2. 2.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
@@ 2 Informal Overview
In this section, we will give a high-level overview of the main ideas and the algorithm. We first introduce the overall approach:
@@ 2.1 The overall approach
From now on, unless otherwise noted, we will assume that we have a known upper bound $\Delta_{\max}$ on the maximum degree $\Delta$. This assumption will later be removed by using a degree-scheduler identical to the one used by Duan, He, and Zhang [DHZ19].
@@
We will take a two-step approach that, at a high-level, is inspired by the distributed algorithm of Ghaffari, Kuhn, Maus, and Uitto [GKMU18]. The idea is to partition the graph into a set of at most $\tfrac{\Delta}{t}$ subgraphs $\mathcal{G}_{h}$ such that $\{E(G_{i})\}_{G_{i}\in\mathcal{G}_{h}}$ partitions $E(G)$ and such that $\Delta(G_{i})\leq(1+\tfrac{\varepsilon}{2})\cdot{}t$ for all $G_{i}\in\mathcal{G}_{h}$. Then one can colour each $G_{i}\in\mathcal{G}_{h}$ with a separate palette using some algorithm that is efficient in low-degree graphs, and then combine all of these colourings into a $(1+\varepsilon)\Delta$-edge-colouring of the original graph $G$.
@@
As a warm-up, we briefly sketch how to boot-strap a simple static algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring in near-linear time using the algorithms of Gabow et Al. [${}^{+}$85]. This algorithm was recently, and independently, obtained by Elkin and Khuzman [EK24], while this manuscript was under preparation, and we therefore refer to their manuscript for a formal description of the algorithm.
@@
The idea is the following: a degree-splitter with discrepancy $\delta$ is a partition of the edge set into two sets $E_{1}$ and $E_{2}$, such that for all vertices $v$, we have that $d_{G[E_{1}]}(v),d_{G[E_{w}]}(v)\leq\tfrac{\Delta}{2}+\delta$, where $G[E_{i}]$ is the graph induced in $G$ by the set $E_{i}$ i.e. the subgraph of $G$ containing all edges of $E_{i}$. This is a relaxed definition of degree-splitting similar to the one used in [GS17], but other variants have also been studied (see for example [${}^{+}$20]). Given an efficient degree-splitter with small enough discrepancy, one can easily compute a partition $\mathcal{G}_{h}$ of the desired form by recursively applying the degree splitter to $G_{1}$ and $G_{2}$ to build a hierarchy whose lowest level form $\mathcal{G}_{h}$.
@@
Gabow et Al. [${}^{+}$85] showed how to compute a degree-splitter with a discrepancy of only $2$ in the static setting. By recursively applying this degree splitter roughly $h=\log\Delta-\log\tfrac{\log n}{\varepsilon}$ times (or 0 times if $\Delta=O(\tfrac{\log n}{\varepsilon})$) one can achieve a split into roughly $\tfrac{\Delta}{\tfrac{\log n}{\varepsilon}}$ subgraphs, each with maximum degree at most $\tfrac{\log n}{\varepsilon}+O(\log n+\sum_{i=0}^{\infty}\tfrac{2}{2^{i}})$. By rescaling $\varepsilon$ a bit, this yields the desired partition. Finally, one can colour each graph in $\mathcal{G}_{h}$ with a distinct palette efficiently using another algorithm of Gabow et Al. [${}^{+}$85].
@@
As we have already briefly touched upon in section 1.1, the two main challenges with making such an approach dynamic are the following: 1)1)1 ) how to maintain the degree splitters, and 2)2)2 ) how to efficiently colour the low-degree graphs.
@@
As for 1)1)1 ), we note that first of all, as far as we are aware, no deterministic dynamic degree-splitting algorithm is known. While it is easy to devise a randomised such algorithm against an oblivious adversary for graphs with not too small maximum degree (simply partition the edges u.a.r into $2$-edge sets), it is unclear how to do this deterministically. Second of all, if a deterministic degree-splitter has recourse at least $10$, then naively maintaining a hierarchy of degree-splitters becomes infeasible as the recourse at depth $h$ of the hierarchy can be as large as $10^{h}$. In order to split the degree from $n^{0.75}$ to $n^{0.25}$, we would require $h\geq\tfrac{1}{2}\log n$, thus resulting in super-linear recourse, and therefore a super-linear update time.
@@
We show in Section 4 that we can, in fact, maintain a dynamic $t$-splitter in $\tilde{O}(\operatorname{\text{{\rm poly}}}(t,\log(n),\varepsilon))$ update time, but crucially with only $\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon)$ recourse. This data-structure is far from trivial, and requires many ideas of its own. We will discuss it in Section 2.2.
@@
$R(h)\leq\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon)^{h}=2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$

### v2

@@ Abstract
Given a dynamic graph $G$ with $n$ vertices and $m$ edges subject to insertions and deletions of edges, we show how to maintain a $(1+\varepsilon)\Delta$-edge-colouring of $G$ without the use of randomisation. More specifically, we show a deterministic dynamic algorithm with an amortised update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ using $(1+\varepsilon)\Delta$ colours. If $\varepsilon^{-1}\in 2^{O(\log^{0.49}n)}$, then our update time is sub-polynomial in $n$.
While there exists randomised algorithms maintaining colourings with the same number of colours [Bhattacharya, Costa, Panski, Solomon SODA’24, Christiansen STOC’23, Duan, He, Zhang SODA’19] in polylogarithmic and even constant update time, this is the first efficient deterministic algorithm to go below the greedy threshold of $2\Delta-1$ colours for all input graphs. On the way to our main result, we show how to dynamically maintain a shallow hierarchy of degree-splitters with both recourse and update time in $n^{o(1)}$. We believe that this algorithm might be of independent interest.
@@ 1 Introduction & Related Work
Edge-colouring is a classic and fundamental problem studied extensively in both graph theory as well as many different models of computation [Gabow, 2BERNSHTEYN, Christiansen, duan, DBLP:conf/soda/BhattacharyaGW21]. Given a simple graph $G=(V,E)$, a $k$-edge-colouring of $G$ is a function $c:E(G)\rightarrow[k]$ that maps every edge to a colour in some colour set $k]=\{1,2,\dots,k\}$, An edge-colouring is proper if $e\cap e^{\prime}\neq\emptyset$ implies that $c(e)\neq c(e^{\prime})$ for any pair of distinct edges $e,e^{\prime}$. The chromatic index $\chi^{\prime}(G)$ of $G$ is then the smallest $k$ for which a proper $k$-edge-colouring of $G$ exists. It is not difficult to check that $\chi^{\prime}(G)\geq\Delta$ if $G$ has maximum degree $\Delta$. A classic result of Vizing [vizing1964estimate] shows that in fact $\chi^{\prime}(G)\leq\Delta+1$. Deciding whether $\chi^{\prime}(G)=\Delta$ or $\chi^{\prime}(G)=\Delta+1$ is, however, not easy, and it was shown by Holyer that determining $\chi^{\prime}(G)$ exactly is NP-complete [Ian].
@@
Recently, the edge-colouring problem has received a lot of attention in the dynamic algorithms community. Here one studies dynamic graphs, where edges are continually inserted and deleted, and the challenge is to maintain some proper edge-colouring of the graph as it is subjected to these updates. There has been some exciting progress on developing efficient randomised dynamic algorithms using only $(1+\varepsilon)\Delta$ colours. This number of colours go significantly below the greedy-threshold of $2\Delta-1$ colours. Duan, He, and Zhang [duan] showed how to maintain such a $(1+\varepsilon)\Delta$-edge-colouring with an update time of $O(\tfrac{\log^{7}n}{\varepsilon^{2}})$ provided that $\Delta=\Omega(\tfrac{\log^{2}n}{\varepsilon^{2}})$. Christiansen got rid of all requirements on $\Delta$ while maintaining a worst-case update time in $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$. Very recently Bhattacharya, Costa, Panski, and Solomon [CostaEtAl] pushed the update time all the way down to $\tilde{O}(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ under some additional assumptions on $\Delta$ and $\varepsilon$.
@@
However, all of the above algorithms rely on randomness in several and crucial ways, and the landscape of deterministic dynamic algorithms is a lot less colourful. Barenboim and Maimon [barenboim2017fully] gave a dynamic algorithm with $O(\sqrt{\Delta})$ worst-case update time using $O(\Delta)$ colours. This was subsequently improved by Bhattacharya, Chakrabarty, Henzinger and Nanongkai [bhattacharya2018dynamic] who gave a fully dynamic algorithm for computing a $(2\Delta-1)$-edge-colouring with worst-case $O(\log\Delta)$ update time. Bhattacharya, Costa, Panski, and Solomon [DBLP:journals/corr/abs-2311-08367] and independently Christiansen, Rotenberg, and Vlieghe [Vlieghe] showed how to maintain a $(\Delta+O(\alpha))$-edge-colouring in amortised $O(\operatorname{\text{{\rm poly}}}(\log n))$ update-time, where $\alpha$ is the arboricity of $G$. However, if $\Delta\sim\alpha$, this algorithm will not always use a number of colours below the greedy threshold as $\alpha$ can be as big as $\tfrac{\Delta}{2}$. In particular, until now no deterministic and dynamic algorithm always using less than $2\Delta-1$ colours with sub-linear update time was known.
@@
Theorem 1.1 (Informal).
Let $G$ be a dynamic graph and $\varepsilon^{-1}\leq 2^{O(\log^{0.49}n)}$ a given parameter. Then, we can maintain a proper $((1+\varepsilon)\Delta)$-edge-colouring of $G$ in $n^{o(1)}$ amortised update time per operation.
@@ 1.1 Main technical challenges
In this section, we will pinpoint two issues towards designing deterministic dynamic edge-colouring algorithms with few colours. These are barriers we know how to overcome in either the static setting or by using randomisation, but so far, to the best of our knowledge, no current deterministic and dynamic techniques can overcome them.
@@
Many static algorithms, distributed algorithms, and dynamic randomised algorithms produce edge-colourings with few colours via a two-step approach. The first step is a form for dimensionality reduction; one attempts to reduce the problem to a case where $\Delta$ is small. The second step is then to apply an algorithm that is fast and efficient for graphs with small maximum degree.
@@
Many static algorithms [Gabow, Costa2, Sinnamon] and many distributed algorithms [ghaffari2018deterministic, split1, split2] address the first step by computing a hierarchy of degree-splitters. A degree-splitter splits a graph $G$ into two covering subgraphs $G_{1},G_{2}$ such that the maximum degree of $G_{i}$ is roughly half of the maximum degree of $G$. By applying this procedure recursively one can reduce the problem to $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ sub-problems each of maximum-degree $\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)$. However, these static approaches for splitting the degree are very precise, and therefore seem prone to very high recourse in the dynamic setting. Hence, we shall take an approach much more resilient to local changes in the graph.
@@
If the maximum degree is large enough, it is easy to realise something similar in the randomised setting, even dynamically against an oblivious adversary. Indeed, one can place each edge into one of $\tfrac{\Delta}{\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n)}$ subgraphs uniformly at random. It then follows from a standard application of Chernoff-bounds that the maximum degree of each such subgraph is split nicely with high probability. As soon as the adversary becomes adaptive, however, this approach immediately breaks down. In the randomised adaptive setting, one can instead rely on a different approach for reducing $\Delta$: one can sample uniformly at random a palette $S\subset[t]$ of colours such that the graph induced by all edges coloured with a colour from $S$, say $G[S]$, exhibits useful properties. Indeed, one can show that with high-probability the maximum degree of $G[S]$ is small and every vertex in $G$ has a free colour in $S$. One can then colour a new edge, by viewing it as being inserted into $G[S]$. This is the approach taken in [Christiansen, duan]. It is, however, not clear how to perform such a reduction in the deterministic dynamic setting; naively building a hierarchy can easily result in an insurmountable recourse (see Section 2), and sampling colours or edges relies crucially on the power of randomisation.
@@
The second challenge is to efficiently colour a low-degree graph. In the static setting one could rely on parallelism [Gabow], and in the dynamic & distributed settings recent progress has shown that the recourse for extending colourings on low-degree graphs is small [2BERNSHTEYN, BernDhawan, Christiansen, duan] and, more importantly, can be efficiently realised using randomisation [2BERNSHTEYN, BernDhawan, Christiansen, duan]. However, naively searching for such an extending sequence of recolourings can be expensive, and so it is unclear how to locate it efficiently without using randomisation.
@@ 1.2 Further Related Work
Gabow, Nishizeki, Kariv, Leven, and Tereda [Gabow] showed how to compute a $(\Delta+1)$-edge-colouring in $\tilde{O}(m\sqrt{n})$ or $\tilde{O}(m\Delta)$ time. This improved over the $O(mn)$ time algorithm implied by the work of Vizing [vizing1964estimate]. Recently this was improved by Sinnamon [Sinnamon], and in special settings by Bernshteyn and Dhawan [BernDhawan] and by Bhattacharya, Costa, Panski, and Solomon [Costa2]. Alon [alon2003simple] and Cole, Ost and Schirra [cole2001edge] gave a near linear-time algorithm for $\Delta$-edge-colouring bipartite graphs. By applying a reduction due to Karloff and Shmoys [karloff1987efficient], these algorithms yield $3\lceil\frac{\Delta}{2}\rceil$-edge-colouring algorithms for general graphs. When $\Delta=\Omega{(\varepsilon^{-1}\log n)}$, Duan, He and Zhang [duan] gave a $O(\varepsilon^{-2}m\log^{6}n)$ time algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring. This has recently been improved by Bhattacharya, Costa, Panski, and Solomon [CostaEtAl] to $\tilde{O}(m\operatorname{\text{{\rm poly}}}(\varepsilon^{-1}))$ time for many values of $\Delta$ and $\varepsilon$. A very recent line of work [New1, New3, New2] has managed to bring the time needed for computing a $(\Delta+1)$-edge-colouring in a static graph all the way down to $O(m\log\Delta)$.
The edge-colouring problem has also been extensively studied in the LOCAL model [2BERNSHTEYN, BernDhawan, pettie, Christiansen, ghaffari2018deterministic, split1, split2, su2019towards] (see for instance [pettie, ghaffari2018deterministic] for extensive surveys). Some of these algorithms either rely on or directly consider the degree-splitting problem [ghaffari2018deterministic, split1, split2]. There has also been a lot of recent progress in online models, see for example [ONew1, ONew2, ONew3].
@@ 1.3 Results
Theorem 1.2.
Let $G$ be a dynamic graph, and $0<\varepsilon<1$ a given parameter. We can maintain a proper $(1+\varepsilon)\Delta$-edge-colouring of $G$ in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ amortised [^1] Here $\tilde{O}_{\log\varepsilon^{-1}}$ hides $\log\log n$ and $\log\varepsilon^{-1}$ factors. update time per operation.
@@
Theorem 1.3.
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ throughout the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-edge-colouring of $G$ such that the maximum degree of $G[i]$ is at most $(1+\varepsilon)\tfrac{\Delta_{\max}}{t}$. The algorithm has an amortised update-time in $O(t^{3}\varepsilon^{-4}\log^{3}m)$ and an amortised recourse in $O(\tfrac{\log n}{\varepsilon^{2}})$.
@@
Theorem 1.4.
Let $G$ be a dynamic graph, and let $1>\varepsilon>0$ and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
@@
  3. 3.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}\leq\operatorname{\text{{\rm poly}}}{\mathopen{}\left(\tfrac{\log n}{\varepsilon}\right)\mathclose{}}$
@@ 2 Informal Overview
In this section, we will give a high-level overview of the main ideas behind the algorithm. We first introduce the overall approach:
@@ 2.1 The overall approach
From now on, unless otherwise specified, we will assume that we have a known upper bound $\Delta_{\max}$ on the maximum degree $\Delta$. This assumption will later be removed by using a degree-scheduler identical to the one used by Duan, He, and Zhang [duan].
@@
We will take a two-step approach that, at a high-level, is inspired by the distributed algorithm of Ghaffari, Kuhn, Maus, and Uitto [ghaffari2018deterministic]. The idea is to partition the graph into a set of at most $\tfrac{\Delta}{t}$ subgraphs $\mathcal{G}_{h}$ such that $\{E(G_{i})\}_{G_{i}\in\mathcal{G}_{h}}$ partitions $E(G)$ and such that $\Delta(G_{i})\leq(1+\tfrac{\varepsilon}{2})\cdot{}t$ for all $G_{i}\in\mathcal{G}_{h}$. Then one can colour each $G_{i}\in\mathcal{G}_{h}$ with a separate palette using some algorithm that is efficient in low-degree graphs, and then combine all of these colourings into a $(1+\varepsilon)\Delta$-edge-colouring of the original graph $G$.
@@
As a warm-up, we briefly sketch how to boot-strap a simple static algorithm for computing a $(1+\varepsilon)\Delta$-edge-colouring in near-linear time using the algorithms of Gabow et Al. [Gabow]. This algorithm was recently, and independently, obtained by Elkin and Khuzman [DBLP:journals/corr/abs-2401-10538], while this manuscript was under preparation, and we therefore refer to their manuscript for a formal description of the algorithm.
@@
The idea is the following: a degree-splitter with discrepancy $\delta$ is a partition of the edge set into two sets $E_{1}$ and $E_{2}$, such that for all vertices $v$, we have that $d_{G[E_{1}]}(v),d_{G[E_{2}]}(v)\leq\tfrac{\Delta}{2}+\delta$, where $G[E_{i}]$ is the graph induced in $G$ by the set $E_{i}$ i.e. the subgraph of $G$ containing all edges of $E_{i}$. This is a relaxed definition of degree-splitting similar to the one used in [split2], but other variants have also been studied (see for example [split1]). Given an efficient degree-splitter with small enough discrepancy, one can easily compute a partition $\mathcal{G}_{h}$ of the desired form by recursively applying the degree splitter to $G_{1}$ and $G_{2}$ to build a hierarchy whose lowest level form $\mathcal{G}_{h}$.
@@
Gabow et Al. [Gabow] showed how to compute a degree-splitter with a discrepancy of only $2$ in the static setting. By recursively applying this degree splitter roughly $h=\log\Delta-\log\tfrac{\log n}{\varepsilon}$ times (or 0 times if $\Delta=O(\tfrac{\log n}{\varepsilon})$) one can achieve a split into roughly $\tfrac{\Delta}{\tfrac{\log n}{\varepsilon}}$ subgraphs, each with maximum degree at most $\tfrac{\log n}{\varepsilon}+O(\log n+\sum_{i=0}^{\infty}\tfrac{2}{2^{i}})$. By rescaling $\varepsilon$ a bit, this yields the desired partition. Finally, one can colour each graph in $\mathcal{G}_{h}$ with a distinct palette efficiently using another algorithm of Gabow et Al. [Gabow].
@@
As we have already briefly touched upon in Section 1.1, the two main challenges with making such an approach dynamic are the following: $1)$ how to maintain the degree splitters, and $2)$ how to efficiently colour the low-degree graphs.
@@
As for $1)$, we note that first of all, as far as we are aware, no deterministic dynamic degree-splitting algorithm is known. While it is easy to devise a randomised such algorithm against an oblivious adversary for graphs with not too small maximum degree (simply partition the edges u.a.r into $2$-edge sets), it is unclear how to do this deterministically. Second of all, if a deterministic degree-splitter has recourse at least $10$, then naively maintaining a hierarchy of degree-splitters becomes infeasible as the recourse at depth $h$ of the hierarchy can be as large as $10^{h}$. In order to split the degree from $n^{0.75}$ to $n^{0.25}$, we would require $h\geq\tfrac{1}{2}\log n$, thus resulting in super-linear recourse, and therefore a super-linear update time.
@@
We show in Section 4 that we can, in fact, maintain a dynamic $t$-splitter in $\tilde{O}(\operatorname{\text{{\rm poly}}}(t,\log(n),\varepsilon^{-1}))$ update time, but crucially with only $\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon^{-1})$ recourse. This data structure is far from trivial, and requires many ideas of its own. We will discuss it in Section 2.2.
@@
$R(h)\leq\operatorname{\text{{\rm poly}}}(\log(n),\varepsilon^{-1})^{h}=2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$